OpenStudy (anonymous):

the sum of the measures of two angles in a right triangle is 90 degrees, and the difference of their measures is 14 degrees. what are the measures of the two angles?

OpenStudy (anonymous):

\[x+x+14=90\] \[2x+14=90\] \[2x=76\] \[x=38\]

OpenStudy (anonymous):

38 and 52

OpenStudy (anonymous):

Or maybe x+y = 90 x-y = 14 Add-> 2x= 104 -> x =52 so y is 90-52 = 38